OpenStudy (anonymous):

Would this be considered a geometric sequence? 5, 30, 180, 1080...

jimthompson5910 (jim_thompson5910):

5 * what number = 30

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

30 times what number = 180 ?

hartnn (hartnn):

by inspection, u can say its not an arithmetic series. for geometric series, u need to check whether there is common ratio..

OpenStudy (anonymous):

60

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

60*30 is 1800 not 180

OpenStudy (anonymous):

lol sorry i meant 6

jimthompson5910 (jim_thompson5910):

and finally 180 times ____ = 1080 what goes in the blank

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

you're multiplying each term by the same number 6 to get the next term

jimthompson5910 (jim_thompson5910):

so this is definitely a geometric sequence

jimthompson5910 (jim_thompson5910):

arithmetic: you add a constant term to each term to get the next term geometric: you multiply a constant term with each term to get the next term

jimthompson5910 (jim_thompson5910):

ex: 1, 5, 9, 13, 17 is arithmetic because 1+4 = 5 5+4 = 9 9+4 = 13 13+4 = 17

jimthompson5910 (jim_thompson5910):

ex: 3, 6, 12, 24, 48 is geometric because 3*2 = 6 6*2 = 12 12*2 = 24 24*2 = 48